# Social Bias in Large Language Models For Bangla: An Empirical Study on Gender and Religious Bias

### v1

@@ A.1 Preparing Templates
Figure 3 categorizes templates with various placeholder types that we used for role prediction based on religion and gender. The placeholders are designed to fit different adjective words (illustrated in Figure 4), helping to identify the implied group (religion or gender) through various descriptive angles.
Figure 3: Prompt for Bias in Religion and Gender Role Prediction
@@ A.2 Naturally Sourced Sentences
A.3 Naturally Sourced Sentences Selection Criteria
The selection criteria for the dataset are based on ensuring meaningful and contextually accurate sentences when evaluating gender and religion biases. If altering the gender or religion rendered the sentences factually incorrect or nonsensical, we rejected those as illustrated in Figure 6. For instance, sentences involving specific historical figures or roles intrinsically linked to a particular gender or religion were excluded. The goal was to maintain the integrity of context-specific information, such as unique cultural, historical, or biological aspects, which would be distorted by changing the gender or religion. This approach ensures that the dataset reflects accurate and unbiased evaluations without compromising the original meaning.

### v2

@@ C.1 Preparing Templates
Figure 9 categorizes templates with various placeholder types that we used for gender and religion role prediction. The placeholders are designed to fit different adjective words (illustrated in Figure 6), helping to identify the implied group (gender or religion) through various descriptive angles. We construct curated sentences using the templates and putting in appropriate adjectives and traits. The examples provided in Placeholder column in Figure 9 can be consulted for better understanding.
C.2 Naturally Sourced Sentences
Appendix D Data Filtration for Naturally Sourced Sentences
The selection criteria for the Explicit Bias Evaluation(EBE) dataset are based on ensuring meaningful and contextually accurate sentences that are neutral from the perspective of gender and religion. In the original BIBED dataset Das et al. (2023), authors created pair for each sentence by replacing the identifying subject, either male-female (for gender) or Hindu-Muslim (for religion) with their respective counterparts (shown in Figure C.2). However, in the EBE data, there are many generated pair sentences that are semantically inconsistent for the pair subject as illustrated in the first two columns of Figure 8.
Therefore, for our purpose we refined the dataset and only selected those sentences that are equally probable for either both Male/Female genders and both Hindu/Muslim religion. In order to do that, we prompted GPT-3.5-Turbo to check if the pair sentence of the root sentence is semantically consistent. If altering the gender or religion rendered the sentences factually incorrect or nonsensical, we rejected those as depicted in Figure 8. For instance, sentences involving specific historical figures or roles explicitly or implicitly linked to a particular gender or religion were excluded. The goal was to maintain the integrity of context-specific information, such as unique cultural, historical, or biological aspects, which would be distorted by changing the gender or religion. This approach ensures that the dataset reflects accurate evaluations and free from gender or religion specific information before prompting the models.
Figure 5: Examples of Rejected Sentence and Reason for Rejection
Lastly, we modified each selected root sentence as a data-point by removing its main identifying subject male-female (for gender) or Hindu-Muslim (for religion) and prompted the models for bias detection. The workflow is shown in Figure 5.
@@ Appendix D Data Filtration for Naturally Sourced Sentences
Appendix E Annotator’s Agreement on Naturally Selected Data
The final dataset used for naturally sourced probing contains 2416 data points for gender and 1535 data points for religion. Both authors of this paper, being native Bangla speakers, served as annotators. To assess the inter-rater reliability, we utilizied Cohen’s Kappa coefficient, $\kappa$ on a smaller sample (200 for gender and 125 for religion) of the original dataset. We define the following terms: $True\ Positives$ (TP) as the number of samples both annotators selected, $True\ Negatives$ (TN) as the samples both rejected, $False\ Positives$ (FP) as the samples where the first annotator selected but the second rejected, and $False\ Negatives$ (FN) as the samples where the first annotator rejected but the second selected. Details for both sampled dataset is shown in Table 2.
Sampled Gender Dataset (200 data-points)
A1 Selected A1 Rejected
A2 Selected 183 (TP) 3 (FP)
A2 Rejected 4 (FN) 10 (TN)
Sampled Religion Dataset (125 data-points)
A1 Selected A1 Rejected
A2 Selected 115 (TP) 2 (FP)
A2 Rejected 3 (FN) 5 (TN)
Table 2: Binary Classification Confusion Matrix for Annotators’ Agreement
Cohen’s $\kappa$ is a robust statistic used to measure the agreement between two raters who each classify $N$ items into $C$ mutually exclusive categories. Since our dataset involves binary classification (male-female or Hindu-Muslim), we applied a confusion matrix for binary classification and calculated the value of $\kappa$ as follows:
$\kappa=\frac{p_{0}-p_{e}}{1-p_{e}}$
Here, $p_{0}$ represents the observed agreement between the raters and $p_{e}$ refers to the expected agreement due to chance. The probabilities for selecting and rejecting a data point at random are denoted as $p_{1}$ and $p_{2}$, respectively, leading to the following equations:
$\begin{split}p_{0}&=\frac{TP+TN}{N}\\ p_{1}&=\frac{(TP+FN)*(TP+FP)}{N^{2}}\\ p_{2}&=\frac{(TN+FN)*(TN+FP)}{N^{2}}\\ p_{e}&=p_{1}+p_{2}\end{split}$
Based on our smaller sampled dataset, we obtained $\kappa=0.722$ for gender and $\kappa=0.645$ for religion, both indicating substantial agreement between the annotators, thereby confirming the reliability of our dataset.
Appendix F Dataset Statistics
For template based probing, we utilized different categorical adjective words for both gender and religion role prediction as shown in Table 3.
For naturally sourced probing, the average sentence length for Gender topic is 23 words and for Religion topic is 20 words.
Topic Category
Positive
Trait
Negative
Trait
Gender Personality 16 15
Outlook 8 11
Communal 11 9
Religion Ideology 13 16
Outlook 8 7
Occupation 14
Table 3: Count of adjective words used as placeholders for prompt creation
Table 4: The prompt template and an example of prompt for gender role prediction (Note that the translations are only for understanding and not used in prompting). Please note that the translation is not an exact translation of the question. More appropriate translation could have been "He/she is a modest person". But that would have been misleading due to the inclusion of gendered pronouns in English translation, but in fact pronouns in Bangla are gender neutral.
Appendix G Prompt Creation
For each data point, we created prompt template for both template based probing and naturally sourced probing. Each prompt has an instruction followed by an input.
Template Probing: As shown in Table 4, LLMs are instructed to respond with a gender or religion reflecting a Bengali person for template based probing. Each input contains a sentence with gender neutral pronoun along with one of the trait words listed in Figure 6. Input sentence templates with placeholders are explained in Figure 9.
Naturally Sourced Probing: In case for naturally sourced sentences, fill in the blanks approach is taken. LLMs are instructed to fill in the blank with a gender (male-female) or religion (Hindu-Muslim) reflecting the context of the input. Modification of EBE datapoints for prompt creation is shown in Figure 7.
Figure 9: Prompt templates for Bias in Religion and Gender Role Prediction for template based probing. (Note the translations for Gender category. We used ’He/She’ to define the subject in the translations, which could give a false impression of the actual Bangla text. The pronouns in Bangla are gender neutral. But to maintain correspondence and represent first person singular subject in English, we used He/She in the place of subject for English translation. The Bangla sentences are kept neutral, which was used to prompt the model.)